OpenStudy (highschoolmom2010):

Consider a box with dimensions 3, 4, and 5. a. Find its surface area. b. Double each dimension and then find the new surface area. c. Find the ratio of the new surface area to the original surface area. d. Repeat parts (a)–(c) for a box with dimensions 6, 9, and 11. e. Make a Conjecture How does doubling the dimensions of a rectangular prism affect the surface area?

OpenStudy (anonymous):

for a, you multiply all 3 dimensions

OpenStudy (highschoolmom2010):

@ganeshie8 @hartnn @phi

ganeshie8 (ganeshie8):

nice activity, start by sketching a rough diagram of box

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

oh you're still not seeing the diagrams here ?

OpenStudy (highschoolmom2010):

nope

ganeshie8 (ganeshie8):

okay, box is a prism. so you need to use prism formula

OpenStudy (highschoolmom2010):

like that

ganeshie8 (ganeshie8):

yes, looks good ! we can use that pic

ganeshie8 (ganeshie8):

a. Find its surface area.

ganeshie8 (ganeshie8):

use SA formula of prism

OpenStudy (highschoolmom2010):

\[SA=LA+2B\]

ganeshie8 (ganeshie8):

Yes, \(LA = ph\) first find out perimeter of base \(p\)

OpenStudy (phi):

looks like you are in good hands

OpenStudy (highschoolmom2010):

p=(3*4)

ganeshie8 (ganeshie8):

thats not perimeter. remember perimeter is length around base

ganeshie8 (ganeshie8):

p = 3 + 4 + 3 + 4

OpenStudy (highschoolmom2010):

p=2(3+4) p=14

OpenStudy (highschoolmom2010):

i always get them mixed up

ganeshie8 (ganeshie8):

happens lol, LA = ph = ?

OpenStudy (highschoolmom2010):

LA=(14*5) LA=70